OpenStudy (donaldtrump):

What is the product of 5\8 and 3\4? @thesmartone @confluxepic

TheSmartOne (thesmartone):

What is 5 x 3 = ? And What is 8 x 4 = ?

OpenStudy (confluxepic):

Do you know how to multiply fractions? @DonaldTrump @~

OpenStudy (alexandervonhumboldt2):

\[\frac{ 5 }{ 8 }*\frac{ 3 }{ 4 }=\frac{ 5*3 }{ 8*4 }=simplify\]